When 3 times a certain number N is added to 6, the sum is 20 more than the original number. What is the number N?
Answer by kdr(11)

The equation would be:
3N + 6 = N + 20
Subtract N from each side:
2N + 6 = 20
Subtract 6 from each side:
2N = 14
Divide each side by 2:
N =7
Check:
3(7) + 6 = 7 + 20
21 + 6 = 27
27 = 27